OpenStudy (anonymous):

find the general solution of equation xdy-( sqrt(y^2+x^2)+y)dx=0

OpenStudy (zehanz):

WolframAlpha has a well-explained solution:

OpenStudy (anonymous):

thanks alot but i don't understand how do you get(sinh) in the answer;

OpenStudy (anonymous):

xdy -ydx = sqrt(x^2+y^2)dy -(-xdy +ydx)/y^2 = sqrt(x^2+y^2) /y^2 dy or -d(x/y)= sqrt(1+(x/y)^2) /y dy or -d(x/y)/ sqrt(1+(x/y)^2) =1/y dy or ln c -ln(x/y +sqrt(1+(x/y)^2)) =ln y and so on...

OpenStudy (zehanz):

Read up until the step\[\int\limits \frac{ v'(x) dx}{ \sqrt{(v(x))^2+1} }=\int\limits \frac{ 1 }{ x }dx\]Now \[(\arcsinh(x))' =\frac{ 1 }{ \sqrt{x^2+1} }\]So with the Chain Rule (in reverse) the integrals become\[\arcsinh(v(x))=\ln|x| + c\]And\[v(x)=\sinh(\ln|x| +c)\]But: xv(x)=y(x), and this leads to:\[y(x)=x \sinh(\ln|x|+c)\]

OpenStudy (anonymous):

thanks; I understood

OpenStudy (anonymous):

I was trying to solving it cheek this answer please; let x=vy,dx=vdy+ydv vydy-(sqrt(v^2 y^2)+y)(vdy+ydv)=0 -vy sqrt(v^+1)dy-y^2(sqrt(v^2+1)+1)dv=0 -y(vsqrt(v^2+1)dv+y(sqrt(v^2+1)dv))dv=0 let y not equal zero by intergration -ln|y|=ln|v|-arc csch(v)+c

OpenStudy (zehanz):

I think you've forgot to put y² in the second line. It should be: vydy-(sqrt(v²y²+y²)+y)(vdy+ydv)=0

OpenStudy (zehanz):

Then: vydy-(ysqrt(v²y²+1)+y)(vdy+ydv)=0

OpenStudy (anonymous):

oh my god, that's right Sorry for your trouble with me

OpenStudy (anonymous):

thanks again

OpenStudy (abb0t):

variation of parameters.

OpenStudy (anonymous):

what do you mean?

OpenStudy (abb0t):

Are you using variation of parameters to solve? If you don't know what I am talking about then just ignore me. Lol. I don't want to confuse you more.

OpenStudy (anonymous):

I didn't use variation of parameter to solve this equation thanks for your time